How many liters of a 60% alcohol solution must be added to 8 liters of a 10% alcohol solution to get a 44% alcohol solution?
Answer by Alan3354(69443)

How many liters of a 60% alcohol solution must be added to 8 liters of a 10% alcohol solution to get a 44% alcohol solution?
-------------------
Add L liters
8*0.1 + L*0.6 = (8+L)*0.44
8 + 6L = 35.2 + 4.4L
1.6L = 27.2
L = 17 liters